Reports 1-1 of 1 Clear search Modify search
MIF (General)
takafumi.ushiba - 0:04 Wednesday 02 November 2022 (22764) Print this report
PRMI lock trial Day4: test of three-mirror ADSs for PRMI

Continued work of klog22749.

I succeeded to engage ADS for PRMI to three mirrors: IMMT1, IMMT2, and PRM.
However, IMMT1 and IMMT2 moves a lot from the values for FPMI, and IMMT1 actuators were saturated.
So, we need to offload itby usign PICO motors at some point.

First, I repeated yesterday's ADS scheme to confirm the reproducibility and then found that IMMT1 YAW and IMMT2 YAW has a same excitation frequencies.
So, I changed the frequency of IMMT1 excitation to 22.1 Hz to 28.1 Hz.
Also, I changed the PD signals for IMMT1 feedback from POP DC to POP90 because POP90 is more sensitive to IMMT1 motion.

After above changes, ADS seems working properly and POP90I signals became about 1 (fig1), while the recoard yesterday was 0.6.
However, the new good values of IMMTs were far from the good values for FPMI and IMMT1 actuators were saturated during ADS.
So, we need to offload it at some point.
Figure 2 shows the setting of ADS for PRMI today.

Obtained good OpLev values for PRMI are as follows:
IMMT1: (P,Y)=(122,234)
IMMT2: (P,Y)=(-83,120)
PRM: (P,Y)=(-20,-14)

I will implement the new initial alignment scheme for PRFPMI in the guardian soon.

After the work, I restored the good OpLev values of IMMT1 and IMMT2 to the one for FPMI to kep the good alignment for it.

Images attached to this report
Comments to this report:
tomotada.akutsu - 0:41 Wednesday 02 November 2022 (22765) Print this report

Just for my curiosity, to compare signals when POP90I was building up around 1. Although IMMT1 and 2 good oplev values varied like 100-200 urad, signals at POP_forward QPDs seems within the QPD range, maybe..

Images attached to this comment
Search Help